Klapperich Lab:Notebook/Lab Meeting Notes/2010/02/25

From OpenWetWare
Jump to navigationJump to search
Owwnotebook icon.png Project name Report.pngMain project page
Resultset previous.pngPrevious entry      Next entryResultset next.png

25 February 2010 Lab Meeting

Alex Presentation.

‡ Announcements

‡ Flu R01:Integration

  • meetings every other M 1-3pm. 705.

† HDA (Lead: Jaephil, Team:Sonali)

  • Patient Samples - dilution experiments (MM/SH/JD)
  • Start planning R01 for Submission on 6/5/10. MM, JD, CMK.
  • HDA mix in dried form - Expecting it from BH.
  • warm start on the way
  • HDA negative water reactions all have primer dimers/non-specific band at 70bp for c diff HDA (MM)
  • 2nd HDA design - "microSPE" + "Flu primer" + "HDA in a cup" + "Detection kit from Biohelix"

† Sample Concentration (Lead: Jane, Team: Jaephil)

  • Air resistance in the air channel is too large.
    • New branched "membrane" chip design #1 in lithography stage.
    • New "tube" chip design in drawing stage. Obtained more porous PTFE tubing samples
  • Samples expected from Fauchet group.

† SPE Column Optimization for DNA/RNA. (Lead: Sonali, Team: Hussam, Jessie, Brendan)

  • Sonali put out a new RNA protocol (correct SPE batter, homemade buffers, less ethanol washing, RNA Secure) before break.
  • Protocol tested with 10ng, 1ng, 0.1ng Stratagene purified human RNA (MM) and 100 uL of 1:64 dilution ATCC MDCK sup Flu A/PR/8/34 (JC)
  • Human Total RNA see % yields at 50% and fraction 1 + fraction 2 = 4-5ng RNA out of 10ng with all changes. With high vol *ethanol wash but other protocol improvements, see 40% yield at all other concentrations.
  • For Flu virus see improved Ct value by 7 cycles in fraction 1 from multiple channels
  • Also detecting 1.4 logs more of virus at 1:64 JC virus prep than before.
  • Switching from lambda phage DNA to TaqMan RNase P detection [update]
  • Hussam Runs chip with RNA (confirm procedure is correct) [update]
  • Jessie will run chips with serial virus dilutions from 1:1 to 1:1,000,000 to re-do previous dilution experimnent with WHO SYBR green assay on chip.
  • Jessie tried 4 2mm channels with 4uL SPE on Black Beauty. Only one sucessful elution with 70uL in two fractions. Got 1E6 copies compared to 8E6 previously.
  • Sonali completed 5 total patients (BUMC and BIDMC) extracted and RT-PCR'd successfully using new RNA protocol and small SPE channel chips. All samples selected based on Qiagen kit extractions having low Cts at 15-20 cycles
  • Sonali gave Qingqing 3 of these patient RNAs extracted on chip SPE for testing on-chip RT-PCR with SuperScript vs Qiagen one-Step kit.

† Integrated chip for flu(SPE+RT+PCR) (Lead: Qingqing)

  • integrated chip works with ATCC flu virus
  • working on improving the efficiency
  • testing real patient sample
    • talk about PCR improvements BEFORE we use new patient samples.
  • PCR1 will combine with PCR2 if PCR2 comes back for revisions. (CMK)

‡C. diff Project (Cathie, Sonali, Satish Singh, Lisa J., His post doc )

  • See email notes.
  • See HDA notes above for styrofoam cup/chip work (MM)

† PCR of C.Difficile DNA (Qingqing)

  • ATCC C.dif DNA with Negative patient sample works on chip
  • Another positive patient sample was tested on chip
 * wrong size products(55bp).
 * no designed size product.

‡ Coulter Flu Fraunhofer Project (Lead: Sonali, Team: Sonali, Jessie, Cathie, CMI Folks, Qingqing)

  • New meeting time, T, 9:30 am every other week. MEETINGS CANCELLED THIS WEEK.
  • Will be done by next Tuesday - PCR with 3 different kits of same serial dilution RNA sample (Invitrogen - SuperscriptIII Platinum, Ambion - AgPath-ID, Qiagen - OneStep RT-PCR Kit) -- DONE 2/25 br>

‡ Agilent Automated Sample Preparation (Lead: Alex)

  • Paper on HotDog - with CMK/ASB

‡ COBRA (Lead: Jaephil, Team: Cathie, Jane for virus only, PHO folks)

  • Evap paper - Additional experiment for Fig 3. By 3/15.

‡ Biointerfaces group (Lead: MinCheol, Team: Cathie, MCK, Wong and Meller folks)
* Cathie will submit paper.

‡ CIMIT- Sepsis (Lead:Cathie, Team:TBA)

  • Sample collection ongoing.

‡ PATH Grant (Lead:Cathie, Team:Frank, Sean, Mark, Jake Trueb (ME))

  • Submitted Gantt Document to PATH 11/25.
  • Frank: One zero-time 6-straw set run with 5-25% efficiency. Second set in progress (24 Feb).
  • Sean: Started making straws for the new experiments we talked about (DNA weights of 1ng, 10ng, and 100ng through 700nm at SPE 100 monolith at 0-day time point). Testing has been very slow-going due to an excess amount of other school work/tests.
  • Mark: as of 25 Feb. has loaded 6 and 4 month trials with 10ng/straw. Results of the 1ng/straw 0-day trial inconclusive, redo possible. Sticking with 10ng/straw for remaining tests. Loaded 2 month trial on Tuesday 2/2. 1-week trial eluted Monday 2/15 again inconclusive - most likely there was a problem with the straws, a bad lot (same lot as the 1ng/straw test). To redo 1-week trial in the future with new lot of straws. Also troubleshooting with smaller wash volumes (down to 50uL) with results on 2/25. Straw fabrication ongoing.

‡ IIH Senior Project.

  • Progress Report 1 completed